### v1

@@ Abstract
The neuronal electric operation can entirely be described by physics provided that we take into account that slow positive ions instead of fast light electrons are the charge carriers that biology uses. Based on the slow ions, we derive the correct physical model that Hodgkin and Huxley wanted. By using laws of electricity and thermodynamics in a non-disciplinary way, we derive the resting potential from the first principles of science, from the biological "construction" (the thickness of the isolating semipermeable membrane and the presence of ion channels in the walls of the membrane separating ionic solutions). We show that a simple control circuit regulates neuronal operation, between a resting and a transient state and that issuing an Action Potential is part of the control process. We explain that the resting potential is the consequence of charge separation instead of an ad-hoc linear combination of mobilities or reversal potentials as the Goldman-Hodgkin-Katz equation claimed. We defy the hypothesis suggested by Hodgkin and Huxley (HH) that some "leakage current" produces the resting potential; this way solve the decades-old mystery of "heat absorption" by neurons, the controversy of the experienced and theoretically expected energy consumption of neural computing. We fix the mistakes HH guessed in the physical picture behind their empirical model and mathematical equations, and introduce the correct physical model by providing also its detailed mathematical description.
@@
slow ions, neuron’s physical model, resting potential, leakage current, negative feedback, thermodynamic driving force, Goldman-Hodgkin-Katz equation, irreversible heat production
@@ 1 Introduction
The excellent textbook [1] separates the neuron membrane’s states into resting and transient states. Concerning the resting state, the statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; here we tell the second half. We refine their statement "by discussing how resting ion channels establish and maintain the resting potential"  [1], page 126. In the resting state, two effects are considered. On the one side, the potential is created by a static charge-up: "the cell membrane potential results from the separation of net positive and net negative charges on either side of the membrane". Furthermore, as we discussed, it is only a small part of the potential: the polarization of the electrolyte’s dipoles contributes to the rest. (For the sake of clarity: both mechanisms are known in physics. The word ’polarization" is reserved for rearranging charges in multiply-charged systems without external current. The changes in potential difference is the consequence of complete charge separation, even if physiology calls it, for example, hyper-polarization and repolarization. External currents are associated to the changes of potential.) On the other hand, it was claimed that "the resting membrane potential is the result of the [low intensity] passive flux of individual ion species through several classes of resting channels" [1], although it was question marked whether the membrane permeability is a primary property for the membrane potential generation, for a review, see [2]. The claim is rooted in the hypothesis and the measurements by Hodgkin and Huxley [3], as we discuss in section 3.5. Its numeric figure was correct, but it was mistakenly interpreted as being produced by a "leakage current" on the membrane’s parallel resistance. Paper [4] stated correctly "that separation of charges produces a potential" and derived the thermodynamic part in two different ways but did not account for the electric potential. We derive how that electric potential is produced by separating charges and how its value can be calculated from the first principles.
@@
Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water." [5]. More precisely, inseparable thermodynamic and electrical processes keep the potential set by the inseparable electric and thermodynamic backbone defined by the lipid structure of the cell. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [6].
The book asks the central questions, "How do ionic gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, as it discusses only membrane permeability without explaining how the resting potential is created and regulated through charging and polarization.
@@
We demonstrate that classical electric methods enable us to calculate the membrane’s potential resulting from charge separation and polarization. However, some physically plausible assumptions must be made to provide numerical figures. The electric processes are inseparable from the corresponding thermodynamical processes [7], which we discuss by deriving a "thermodynamical electrical field". The other minor effect is a low-intensity "passive flux" (material transport) produced by the "resting channels". They are ion channels not necessarily only across the membrane; they also include the ion channels in the Axon Initial Segment (AIS) [8, 9, 10, 11, 12]. The "resting current" is about two orders of magnitude lower [13] than it was assumed at the time of writing the book; correspondingly, its role is much smaller in the transient state, producing an Action Potential (AP) which we discuss elsewhere [14].
The importance of the subject was correctly estimated decades ago [15]: "For the past forty years our understanding and our methods of studying the biophysics of excitable membranes have been significantly influenced by the landmark work of Hodgkin and Huxley. Their work has had far reaching impact on many different life science sub-disciplines where concepts of cell biology have come to be important. These include not only neurophysiology but also endocrinology, muscle and cardiac physiology, and developmental biology. The ionic currents and electrical signals generated by neuronal membranes are of obvious importance in the nervous system. But ionic fluxes also play important roles in affecting cellular functions such as secretion, contraction, migration, etc." To fix the flawed hypothesis has a similar importance.
@@ 2 Membrane’s electricity
It was already guessed that the membrane’s resting potential is of electrostatic origin [16, 17, 18, 2, 19]. When ions are confined in a closed volume, they exist in a state of thermal and electrical equilibrium. We must discuss them in a non-disciplinary way because the participating ions belong simultaneously to the disciplines of electricity and thermodynamics; furthermore, we must cross the borderline between particle-based and continuous points of view of nature [7]. Furthermore, the features of living matter may change when measured in physics laboratories as usual. When measuring them, we must consider that "the construction is different from anything we have yet tested in the physical laboratory" [20].
In the absence of external influences (that includes the lack of a separating membrane), both gradients are balanced and are at zero. When this state is perturbed, the system attempts to find a new balanced state, using temporal processes. In this scenario, the ’carrier’ - the ion - can be influenced by two interactions. We discussed elsewhere that those phenomena in living matter need deriving and applying ’non-ordinary’ laws of science [7], and the time course of those processes can be described by the laws of motion of biology. Fundamentally, due to the mixing of interaction speeds, considering only one of the interactions leads to incorrect conclusions.
@@ 2.1 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and thermodynamic environments. A molecule has internal electric forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electric field (which can be of electrical or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electric field fluctuates, so their state is dynamic: the molecules dissociate, and the free ions recombine. Depending on their environment, the substances can exist in base, ionized, and polarized states (and behaviors).
In biological cells, a small (approximately $10^{-3}$) portion of the molecules dissociates (i.e., the ions separate from their counterparts with opposite charges), and they move freely within the volume. In other words, the electrolyte liquid can then conduct electricity due to the mobility of the positive and negative ions. The rest of the molecules can be more or less polarized, allowing for the possibility of producing an internal electric field in the solution.
2.2 Deriving "thermal electric field"
In a segmented electrolyte, the electric charges are typically globally balanced; however, they may become unbalanced locally due to physical reasons. The Nernst-Planck electrodiffussion equation in a balanced state
@@ 2.2 Deriving "thermal electric field"
known as Nernst equation. In other words, Eq.(1) and Eq.(2) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration). Hence, the derived potential inherently comprises an additive term (a potential difference), so they are not intercomparable directly if they use a different reference $C_{k}^{int}$. The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [7].
(It is a common fallacy in physiology that the applied potential $U$ generates a current. Instead, its gradient ($E=U/d$) has the effect, where $d$ is the distance, for example, between the clamping point and the membrane or the AIS.)
We assume that, in a balanced state, the concentrations on the two sides of the membrane are $C_{k}^{ext}$ and $C_{k}^{in}$, respectively; furthermore we assume that
@@
To simplify the math writing, we neglect the change caused by considering the difference (maybe we should apply a near-unity factor); we use the larger concentration magnitude instead. Using the constants and $T=300$, and that $U=E*d$, we can derive the equivalent "thermodynamic electric field" for the case of a permeable ion channel in the membrane
@@
It does not depend on the concentrations and depends reciprocally on the membrane thickness (given in $nm$); furthermore, it is valid separately for all $C_{k}$
@@
For a $5\ nm$ membrane thickness, it results in $5.17*10^{6}\ \frac{V}{m}$, which compares well to the value provided by Equ.(18). Figure 1 depicts its dependence on the membrane’s width and how it compares to the charge-generated electric field. Correspondingly,
@@
In simple words, the Nernst-Planck equation states that the changes in concentrations of ions create changes in the electric field (and vice versa), and in a stationary state, they remain unchanged. Notice that in the case of an ion mixture, a joint charge-generated electric field exists, and they equal, per ion, the concentration-generated "electric fields", see section 3.2.1.
@@
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added an external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid.
@@
That means when describing an ionic transfer process, we must not separate the electric current from the mass transfer: they happen simultaneously and mutually trigger each other. Notice that the thermodynamic term is ion-specific while the electric term is not. To be entirely balanced, the system must be balanced to all elements. In this way, changing one concentration implicitly changes all other concentrations and the electric field.
It is a colossal mistake to introduce equivalent circuits with their fixed-value voltage generators. It forces one to assume that the conductances of the players (membrane, synapses, AIS) change without any reason and prevents understanding how the competition between thermodynamic and electric processes governs neuronal operation. It leads, among others, to attributing conductance change to membranes, which are simple isolators with no charge carriers that can implement charge transfer, see section 2.6; in this way attributing the change of an electric entity to the biological material. This assumption neglects the driving force the mentioned forces provide. Instead, it attributes the magic ability to the ion channels that they can change their transmission ability as the actual situation requires.
@@
When discussing balanced states, no transport occurs, so the transport force cancels, and the mobility has no role. The counterforce adapts to the situation. That force may be mechanical: the ions sitting on the surface of the membrane press the surface due to the attractive force of ions, and the membrane mechanically provides the needed counterforce. If the boundary of the segments is not freely penetrable, the counterforce equals the difference between those two forces. In this way, no force acts on the ions; the two gradients persist. When the ions can move freely between the segments, they will move until they produce a concentration gradient (a thermodynamical force) for the given ion that counterbalances the electrical gradient (the electrical force) and then the transport stops. A transport force is needed to reach the Stokes-Einstein speed (see Eq.(27)) in a viscous fluid.
@@
We must not forget that initially, the Nernst-Planck equation described a transfer process (i.e., a dynamic equation), assuming the same speed for mass and charge transfers. We used it only to describe the balanced state (i.e., a static equation at zero speed) and added the term describing external (such as mechanical constraint) force that may be involved; it affects the process, but it does not belong to either of the respective fields. Commonly known mechanical constraints appear in a non-conducting layer called the membrane, where that counterforce prevents ions from penetrating the membrane layers. A membrane is a perfect isolator, i.e., no charge carriers exist between its two surfaces. (There may be ion channels that deliver ions built into the membrane, but it is a different subject.) The charge-up results from charge separation, a mechanism distinct from polarization, as typically described in most biology textbooks. In the form we derived, the equation correctly describes the ion transfer, a current conveying ions with the speed controlled by the actual gradients, instead of assuming an ion current traveling with the apparent speed of electric current in solids.

### v2

@@ Abstract
Physics perfectly describes neuronal operation, provided that we take into account that biology uses slow positive ions instead of electrons as charge carriers and remove untested ad-hoc hypotheses that contradict science’s first principles. We also incorporate recent experimental discoveries into the outdated classic theoretical description. Lipid mechanisms are really very important for cellular biology, but they are certainly not suitable for describing the phenomena we discuss. We introduce the correct physical model, significantly enhancing the classic Hodgkin and Huxley (HH) model; furthermore, the fundamentally bio-electrically triggered operation causes changes in electrical, mechanical, and thermodynamic properties of living matter. We derive the resting potential from first principles of science, showing that it has no relation to an ad-hoc linear combination of mobilities or reversal potentials as the Goldman-Hodgkin-Katz (GHK) equation claimed. Furthermore, we derive an "equivalent thermodynamic electric field" that enables discussion of, among others, the operation of ion channels, their ion selectivity, and voltage sensing. We demonstrate that a simple electrical-thermodynamic control circuit regulates neuronal operation, setting and maintaining a stable resting potential and handling an unstable transient process known as the Action Potential (AP). Its setpoint entirely defines the resting potential, explaining its robustness during growth and evolution. Our cross-disciplinary approach naturally fuses the electrical and mechanical/thermodynamic description of neuronal operation, resolves the decades-old mystery of "heat absorption" and "leakage current" (with their far-reaching consequences), and derives the thermodynamic description of neural computing. We defy that science cannot describe life.
@@
slow currents, neuron’s physical model, resting potential, leakage current, negative feedback, thermodynamic driving force, Goldman-Hodgkin-Katz equation, reversible heat production, thermodynamics of neurons
The table of contents is present only for the comfort of reviewers
@@ 1 Introduction
The excellent textbook [1] separates the neuron membrane’s states into resting and transient states. The statement that "the resting membrane potential results from the separation of charge across the cell membrane" is only half the truth; here we tell the second half. We refine their statement "by discussing how resting ion channels establish and maintain the resting potential"  [1], page 126. In the resting state, two effects are considered. On the one side, the potential is created by a static charge-up: "the cell membrane potential results from the separation of net positive and net negative charges on either side of the membrane". However, as we discuss, it is only a small part of the potential: the polarization of the electrolyte’s dipoles contributes the rest. (For the sake of clarity, both mechanisms are known in physics. The word ’polarization’ is reserved for rearranging charges in multiply-charged systems without the application of an external current. The changes in the membrane’s potential difference are a consequence of complete charge separation (and the movement of slow charges), even if physiology refers to it as polarization. External currents contribute to the changes of potential.) On the other hand, it was claimed that "the resting membrane potential is the result of the [low intensity] passive flux of individual ion species through several classes of resting channels" [1], although it was question-marked whether the membrane permeability is a primary property for the membrane potential generation, for a review, see [2]. The claim is rooted in the hypothesis and the measurements by Hodgkin and Huxley [3], as we discuss in section 2.6.4. Their numeric figure was correct, but they mistakenly interpreted the voltage they measured as being produced by a "leakage current" on the membrane’s parallel resistance. Furthermore, their interpretation leads to causality issues [4]. Paper [5] stated correctly "that separation of charges produces a potential" and derived the thermodynamic part in two different ways, but did not account for the electric potential. We derive how that electric potential is produced by separating charges and how its value can be calculated from first principles.
@@
Our results underpin that "the crucial system in biology isn’t a molecule or a molecular class whatsoever, but the interface created by biomolecules in water" [6]. More precisely, inseparable thermodynamic and electric processes keep the potential set by the inseparable electric and thermodynamic backbone defined by the lipid structure of the cell. It is one of the frequent cases when one effect implements a functionality (the backbone closely defines the frames of the operation), and the other (in this case, combined thermodynamics and electricity) corrects it when it implements a complex operational (dynamical) functionality: "stimulated phase transitions enable the phase-dependent processes to replace each other … one process to build and the other to correct" [7].
@@
The book [1] asks the central questions, "How do ionic [i.e., electrical and chemical] gradients contribute to the resting membrane potential? What prevents the ionic gradients from dissipating by diffusion of ions across the membrane through the resting channels?" However, it leaves them essentially open by giving only a qualitative answer, as it discusses only membrane permeability without explaining how the resting potential is created and regulated through charging and polarization. We demonstrate in section 3 that classical electric methods enable us to calculate the membrane’s potential resulting from charge separation and polarization. However, one can make some physically plausible assumptions to provide numerical figures. The electric processes are inseparable from the corresponding thermodynamical processes [8], which we discuss by deriving an "equivalent thermodynamical electric field" in section 2.3. The other minor effect is a low-intensity "passive flux" (material transport) produced by the "resting channels". They are ion channels, not necessarily only across the membrane; they also include the ion channels in the Axon Initial Segment (AIS) [9, 10, 11, 12, 13]. The "resting current" is about two orders of magnitude lower [14, 15] than it was assumed at the time of writing the book; correspondingly, its role is much smaller in the transient state, and it even drastically changes the electric model used in producing an AP which we discuss elsewhere [16].
@@ 2 Membrane’s electricity
The importance of the subject was correctly estimated decades ago [17]: "For the past forty years, our understanding and our methods of studying the biophysics of excitable membranes have been significantly influenced by the landmark work of HH. Their work has had a far-reaching impact on many different life science sub-disciplines where concepts of cell biology have come to be important. …The ionic currents and electrical signals generated by neuronal membranes are of obvious importance in the nervous system." To fix the flawed and incomplete hypotheses has a similar importance.
It was already stated that the membrane’s resting potential is, of course, of electrostatic origin [18, 19, 20, 2, 21]. The closed volume of the neurons drastically differs from the infinite space: the charges repel each other, but since they experience the counterforce of the limiting surface (the membrane), a static mechanical pressure and a shock wave are also created, that naturally connects the electric and mechanical effects, as well as the electrical, mechanical, and thermodynamic theories of operation. We must discuss them in a non-disciplinary way because the participating ions belong simultaneously to the disciplines of electricity and thermodynamics; furthermore, we must cross the borderline between particle-based and continuous points of view of nature [8]. Furthermore, the features of living matter may change while being measured in physics laboratories as usual: "the construction is different from anything we have yet tested in the physical laboratory" [22].
In the absence of external influences (which includes the lack of a separating membrane), both gradients are balanced and are at zero. When this state is perturbed, the system attempts to find a new balanced state, using temporal processes. In this scenario, the ’carrier’ - the ion - can be influenced by two interactions. We discussed elsewhere that those phenomena in living matter need deriving and applying ’non-ordinary’ laws of science [8] , and the time course of those processes can be described by the laws of motion of biology. Fundamentally, due to the mixing of interaction speeds, considering only one of the interactions (the disciplinary way) leads to incorrect conclusions and theoretical interpretations.
2.1 Condenser
Biology represents a complex case where the phenomena cannot be reasonably reduced to a single discipline, as classic inanimate science does when deriving its laws for "material points", as A. Einstein coined, abstracted to either mass or charge, but not both. Cell biochemistry sufficiently well describes the lipid bilayers that constitute the membrane [23, 24], but does not explain that the negative and positive ends of lipids in an electrolytic solution attract and absorb ions, forming a charged layer on both surfaces of the membrane. In this way, the membrane in electrolytes acts as a condenser. However, this condenser comprises two charged sheets on the surfaces of the membrane with ill-defined parameters. Their boundaries, thicknesses, and compositions depend on the operating state (as opposed to conducting plates with well-defined sizes and boundaries). This condenser is not ideal: the charged sheet comprises ions with attached aqueous constituents from the electrolytes, so its thickness is far from the idealized zero-thickness known from classical electricity. Furthermore, in its wall, it inseparably comprises ion channels (from an electric point of view: resistors), and the condenser plates are in polarizable electrolyte solutions.
Cell science does not establish its electrical phenomena; instead, it attempts to elucidate complex lipid activities that explain charge transfer, resting and transient potential, and the entire biological operation using mathematical formulas without a realistic physical background. Electrophysiology combines currents of electrons and ions, and uses Ohm’s Law for its admittedly non-Ohmic systems. It introduces foreign (clamping) currents into biological systems, attributing their effect to a changed conductance of the membrane without explaining how the voltage and current can be independent of charge carriers.
@@ 2.2 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and electrical/mechanical/thermodynamic environments. A molecule has internal electric forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electric field (which can be of electrical or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electric field fluctuates, so their state is dynamic: the molecules dissociate, and the free ions recombine. Furthermore, the process plays an important role also in the energy supply of the neurons: the Adenozine Triphosphate (ATP) hydrolyzes molecules and the electric field near the membrane moves the ions to the "condenser plates". Depending on their environment, the substances can exist in base, ionized, and polarized states (and behaviors).
In biological cells, in a resting state, a small (approximately $10^{-3}$) portion of the molecules dissociates (i.e., the ions separate from their counterparts with opposite charges), and they move freely within the volume. In other words, the electrolyte liquid can then conduct electricity due to the mobility of the positive and negative ions (of course, if some electric/thermodynamic field provides the required driving force). The rest of the molecules can be more or less polarized, allowing for the possibility of producing an internal electric field in the solution. In a transition state, all components (including ions, molecules, and ATP) have gradients and they move towards their balanced state.
2.3 Deriving the "thermal electric field"
In a segmented electrolyte, the electric charges are typically globally balanced; however, they may become unbalanced locally due to physical reasons. The Nernst-Planck electrodiffusion equation in a balanced state
@@ 2.3 Deriving the "thermal electric field"
known as Nernst equation. In other words, Eq.(1) and Eq.(2) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration $C_{k}^{int}$). Hence, the derived potential inherently comprises an additive term (a potential difference), so they must not be compared directly if they use a different reference $C_{k}^{int}$ (furthermore, it is nonsense to combine additively quantities from the intracellular and extracellular sides, whether concentrations, or mobilities, or Nernst voltages, as it happens in the GHK equation (see section 5.3). The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [8]. In ’ordinary’ physics, where the charge and mass are independent, changing one side changes the other in the opposite direction. In ’non-ordinary’ physics [26], the two differentials must change in the same direction given that ions’ charge and mass cannot be separated (BTW: this is why Eq. (2) results in a negative potential difference).
It is a common fallacy in physiology that the applied potential $U$ generates a current. Instead, its gradient ($E=U/d$) has the effect, where $d$ is the distance, for example, between the clamping point and the membrane or the AIS. We assume that, in a balanced state, the concentrations on the two sides of the membrane are $C_{k}^{ext}$ and $C_{k}^{in}$, respectively; furthermore we assume that
@@
To simplify the math writing, we neglect the change caused by considering the difference (maybe we should apply a near-unity factor); we use the larger concentration magnitude instead. Using the constants and $T=300$, furthermore, that U=EdU=E*d, we can derive the "equivalent thermodynamical electric field" for the case of a permeable ion channel in the membrane
@@
It does not depend on the concentrations, but depends linearly on the temperature and reciprocally on the membrane thickness (given in $nm$); furthermore, it is valid separately for all $C_{k}$
@@
Figure 1 depicts the field’s dependence on the membrane’s width and how it compares to the charge-generated electric field. The figure illustrates the setpoint, the boundaries of the changes, and that the interesting phenomena happen within the cross-section of the shaded areas. The figure is an illustration, but the figures are not far from the true ones. For a 5nm5\ nm membrane thickness, it results in $5.17*10^{6}\ \frac{V}{m}$, which compares well to the value provided by Eq. (18). (Notice that the membrane’s thickness may change during the action potential, see [27]: the rush-in extra charge compresses the membrane. Similarly, the concentration of the electrolyte in proximity to the membrane also changes.) Correspondingly,
@@
In simple words, the Nernst-Planck equation states that the changes in concentrations of ions create changes in the electric field (and vice versa), and in a stationary state, they remain unchanged. Notice that in the case of an ion mixture, joint charge-generated electric fields exist, and they equal, per ion, the concentration-generated "electric fields", see section 3.1.1.
@@
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added an external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid. (An excellent summary of the possible biology-related forces is given in [28], with a lot of cited references.)
@@
That means when describing an ionic transfer process, we must not separate the electric current from the mass transfer (in other words, to discuss electrical or thermodynamic models): they happen simultaneously and mutually trigger each other. Notice that the thermodynamic term is ion-specific while the electric term is not. To be entirely balanced, the system must be balanced to all elements. This way, changing one concentration implicitly changes all other concentrations and the electric field.
Notice the special role of the constraint force by the closed volume represented by the membrane’s wall. The thermodynamic contribution also means a pressure from the colliding particles, and the electric repulsion exerts a force on the membrane’s wall. Those disciplinary forces act simultaneously and inseparably.
2.4 Fallacies about biological electricity
@@
When discussing balanced states, no transport occurs, so the transport force cancels, and the mobility has no role (when speaking about mobility provided by protein mechanisms, one should derive the corresponding contribution to Eq. (7)). The counterforce adapts to the situation. That force may be mechanical: the ions sitting on the surface of the membrane press the surface due to the attractive force of ions on the opposite side, and the membrane mechanically provides the needed counterforce. This way, no resultant force acts on the ions; the two gradients persist. When the ions can move freely between the segments, they will continue to move until they produce a concentration gradient (a thermodynamic force) for the given ion that counterbalances the electrical gradient (the electrical force), and then the transport stops. A transport force is needed to reach the Stokes-Einstein speed (see Eq. (27)) in a viscous fluid.
@@
We must not forget that initially, the Nernst-Planck equation described a transfer process (i.e., a dynamic equation), assuming the same speed for mass and charge transfers. We used it only to describe the balanced state (i.e., a static equation at zero speed) and added the term describing external (such as mechanical constraint) force that may be involved; it affects the process, but it does not belong to either of the respective fields. Commonly known mechanical constraints appear in a non-conducting layer called the membrane, where that counterforce prevents ions from penetrating the membrane layers. A membrane is a perfect isolator, i.e., no charge carriers exist between its two surfaces. (There may be ion channels that deliver ions built into the membrane, but it is a different subject.) The charge-up results from charge separation, a mechanism distinct from polarization, as typically described in most biology textbooks. In the form we derived, the equation correctly describes ion transfer, a current conveying ions with the speed controlled by the actual gradients, instead of assuming an ion current traveling with the apparent speed of electric current in solids.
It was a colossal mistake to forget that biological currents comprise ions with their electrostatic forces. It led to fallacies, among others, including leakage current (see Section 2.6.4) and heat absorption (see Section 4.6). Their finite (Stokes-Einstein) speed, see section 2.9, due to the ions’ mass and the different conduction mechanism, explains the spatiotemporal behavior of living matter [29] (contrasted to Newton’s ’instant interaction’).
It was another colossal mistake to introduce equivalent circuits with constant-value voltage generators. It forces one to assume that the conductances of the players (membrane, synapses, AIS) change without any reason, and prevents understanding how the competition between thermodynamic and electric processes governs neuronal operation. It leads, among others, to attributing conductance change to membranes, which are simple isolators with no charge carriers that can implement charge transfer, see section 2.9; this way attributing the change of an electric entity to the biological material. This assumption neglects that a driving force is needed to move ions (denies Newton’s laws of motion). Instead, it attributes the magic ability to the ion channels that they can change their transmission ability as the actual situation requires.
The classic textbook [30], appendix A, correctly warned (especially when working with concepts of an electric theory of neuronal operation): "An intuitive grasp of concepts of electricity and electrical circuits is helpful for understanding some of the basic theory in cellular neurophysiology. …Under ideal circumstances, the physical act of measuring a neurophysiological event would have no effect on the electrical signal of interest. Unfortunately, this is seldom the case in neurophysiology."
One of the bad examples is grasping the electrical characteristics by clamping (we are not speaking against the method, but emphasizing the need for its correct interpretation). The textbook also provided the correct interpretation of the measured quantities: "In describing the resistance (or Ohm’s law) properties of neurons, one either injects current and measures the resulting change in voltage (a current clamp) or changes the voltage to different values and measures the current necessary to hold (or clamp) the voltage to these values (a voltage clamp)."
As we discuss in section 4.1, physiology does not realize that in resting and transient states, different physical processes take place and correspondingly, different circuit models shall be used. In the resting state, the low intensity ion current is "shunted" by the ion channels in the membrane’s wall; moreover, the "parallel RCRC" model works sufficiently well. Since the current is slow, it is not simply distributed, but only the "overflow" reaches the AIS. However, as the current intensity exceeds the ion transmission capacity of the "resting ion channels", the "transient ion channels" play an increasing role. Above the voltage threshold, the serial $RC$ circuit is the correct model. Until the neuron’s voltage reaches its threshold, a "parallel" and a "serial" resistor work simultaneously, in this way implementing a mixed model. Measuring the "input resistance" of neurons actually measures, in this sense, a mixture of the "input" and "output" resistance corresponding to the two models, given that the current is distributed between the one through the membrane and the other through the AIS; plus the measuring electrode represents one more drain.
At the beginning, it was clearly told [31], but forgotten: "Since it is quite generally believed that the depolarization of a nerve fiber membrane, during excitation and propagation, involves an increased permeability to ions there have been many attempts to detect and measure this change as an increase in the electrical conductivity. … In these cases, the measuring current was also the stimulating current, and it was not possible to analyze the changes satisfactorily." Moreover, during the passage of an impulse, a "charge wave" passes along the axon and the changed number of charge carriers in that cross-section, see Eq.(20), changes the current that creates the illusion that the conductance at the measured cross section changed. In our terminology, the stabilizing feedback voltage provides an external force, and it also contributes an external current. The feedback amplifier’s electron current (converted to ion current) is provided externally and is added to the closed system, without including the measurement device into the closed system. That interpretation of a closed system denies the fundamental symmetries of science (among others, conservation of charge, mass, impulse, and energy). Moreover, it significantly contributes to the claim that "laws of science cannot describe life".
All conservation laws are correct and valid, only "the construction is different" [22] and the "test method" is wrong. Not mentioning that under some conditions, the "system under test" also "produces" charge carriers by opening ion channels. The "construction" needs revisiting concepts of electricity for living matter instead of projecting concepts valid for inanimate matter only, to living matter. That wrong "grasp of concepts of electricity" suggests that Ohm’s Law is not valid for living systems. The correct statement is that Ohm’s Law is not valid for arbitrarily chosen currents and voltages. Without that revisiting, by saying that living systems show non-Ohmic behavior, biology claims that it applies good laws inappropriately to a good system. The consequences are far-reaching: physiology uses magic changing conductances in circuits with fixed-voltage batteries, that blocks not only understanding the thermodynamic and mechanical effects (and introducing the corresponding theories [32, 33, 34, 35, 36, 27, 37, 38, 6, 39, 40, 41]), but even consistently describing the electric phenomena. We are not speaking against using the concept of conductance, but rather against deriving its value incorrectly by including foreign current compositions, which obscures the actual electrical behavior of biology. In addition, by introducing current feedback (which, by definition, compensates for any gradient in the system), physiology applies an opposite-phase control unit against the neuron’s native control unit and concludes that, after compensation, there is no gradient in the system; in other words, that life exists without needing a driving force.
Physiology is not aware that, inside the neuron, there is a condenser with a voltage difference between its plates. Instead, it believes that a ‘leakage current’ flows through resistors distributed over the membrane. That current is special in that, among other things, it has neither a chemical composition nor charge carriers; furthermore, it can not only emit but also absorb heat. Neuroscience observes the thermodynamic and mechanical changes (for a comprehensive review, see [33]) but cannot connect them to the electrical phenomena. Lacking the required physical background, biology claims that physical laws cannot describe processes in living matter, without providing a plausible and self-consistent theoretical model.